A dart thrower hits the bullseye with his eyes closed 75% of the time. For a bout of throwing that contains 20 tosses, what is the variance for the number of bullseyes he will hit?
Answer by stanbon(75887)

A dart thrower hits the bullseye with his eyes closed 75% of the time. For a bout of throwing that contains 20 tosses, what is the variance for the number of bullseyes he will hit?
----
variance = npq = 20*(3/4)(1/4) = 60/16 = 15/4 = 3 3/4
------------------------------------------
Cheers,
Stan H.
===============